# Two-Level Distributed Interference Management for Large-Scale HAPS-Empowered vHetNets

### v1

@@ II-B HAPS-UE Channel
$\mathbf{h}_{u}^{B+1}=\frac{1}{\sqrt{{PL}_{(B+1),u}}}\Big{(}\sqrt{\frac{1}{1+K_{u}}}\overline{\mathbf{h}}_{u}+\sqrt{\frac{K_{u}}{1+K_{u}}}\overline{\mathbf{h}}_{u}\Big{)},~{}\forall u,\qquad\qquad\qquad$ (3)
where ${PL}_{(B+1),u}$ represents the FSPL between HAPS and UE $u$, calculated according to (2). $K_{u}$ is the Rician factor for UE $u$, and $\overline{\mathbf{h}}_{u}\in\mathbb{C}^{N_{B+1}}$ represents the NLoS component of the channel vector with its elements from a normal random distribution with zero mean and unit variance, $\mathcal{NC}(0,1)$. Accordingly, $\overline{\mathbf{h}}_{u}$ is the LoS component of the channel vector, given by [34] as
$\overline{\mathbf{h}}_{u}=\mathbf{a}(\theta_{u},\phi_{u})\otimes\mathbf{b}(\theta_{u},\phi_{u}),~{}\forall u,$ (4)
@@
Figure 2: HAPS antenna architecture.
As mentioned earlier, we adopt the cell-free as the underlying network architecture where all UEs are served by all BSs. We assume that each BS has perfect information regarding the channel coefficients between antenna elements and UEs. Accordingly, the SINR of UE $u$ in the cell-free vHetNet can be expressed as [35]
$\gamma_{u}=\cfrac{\left|\sum_{b\in\mathcal{B}}\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{u}\right|^{2}}{\sum_{\mathcal{U}\setminus\{u\}}\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}}\right|^{2}+\sigma^{2}_{n}},~{}\forall u,$ (6)
where $\mathbf{w}^{b}_{u}\in\mathbb{C}^{N_{b}}$ refers to the beamforming vector for UE $u$ at BS $b$, and $\sigma^{2}_{n}$ is the variance of the additive white Gaussian noise (AWGN) at $u$. According to the above explanation, the preliminar PFBWD optimization problem can be expressed as
@@
where $\mathbf{W}^{b}=[\mathbf{w}^{b}_{1},\dots,\mathbf{w}^{b}_{U}]\in\mathbb{C}^{N_{b}\times U}$ denotes the beamforming weight matrix at BS $b$, $P_{b}^{\text{max}}$ is the maximum allowable transmit power of BS $b$, and $\gamma_{\text{min}}$ is the minimum required SINR for each UE. In problem (II-B), constraint (II-B) ensures that each UE $u$ meets its minimum data rate requirement, while constraint (II-B) restricts the total transmit power of each BS $b$ to not exceed $P_{b}^{\text{max}}$. Problem (II-B) is nonconvex due to the objective function (II-B) and the nonconvex nature of constraint (II-B), rendering it intractable for conventional solvers. Typically, problems of this nature can be addressed by developing an iterative algorithm using the SCA framework. In each iteration of the SCA, a convex approximation of the original problem is solved. The solution obtained at convergence is considered suboptimal relative to the original nonconvex problem. Additionally, the approximated problems can be solved using either a centralized or a distributed algorithm. However, for large-scale problems, it is advisable to avoid a centralized algorithm, as it requires collecting channel coefficients from all BSs to solve the optimization problem, and thereafter distributing the decisions back to the BSs. This process imposes significant signaling overhead on the network. Additionally, the centralized approach must solve a large-scale optimization problem to determine the decision variables for the entire network, resulting in high computational complexity.
The problem at hand is large in scale and complexity, thereby necessitating the development of a distributed algorithm to solve it efficiently. To enable distributed computation, the original centralized formulation presented in (II-B) must first be transformed into an equivalent distributed formulation to which a distributed optimization method can be applied. The key objective of this reformulation is to enable parallel decomposition, whereby the problem is structured in a way that enforces per-BS constraints. This structure allows the overall problem to be divided and solved across multiple BSs simultaneously. In the subsequent section, we focus on this transformation process, aiming to recast the original problem (II-B) into a distributed format suitable for distributed implementation.
@@ III Distributed Reformulation of the Original Problem (II-B)
In this section, we detail the series of reformulations required to transform the problem (II-B) to a distributed format that can be solved by the two-level distributed algorithm. To this end, first, we deal with the objective function (II-B). We define $U$ slack variables $t_{u},~{}\forall u,$ as the lower bound for $\log_{2}(1+\gamma_{u}),~{}\forall u$. Accordingly, the objective function can be replaced by $f(\mathbf{t})=\sum_{u=1}^{U}\log_{2}(t_{u})$, where $\mathbf{t}=[1,\dots,t_{U}]^{\top}\in{\mathbb{R}^{+}}^{U}$. The $U$ newly defined constraints will be added to the problem as
@@
Next, to deal with the nonconvexity involved in (II-B), we replace it equivalently as
@@
$\beta_{u}\geq\sum_{\mathcal{U}\setminus\{u\}}\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}}\right|^{2}+\sigma^{2}_{n},\,~{}\forall u,$
where $\alpha_{u},~{}\forall u,$ and $\beta_{u},~{}\forall u,$ are new slack variables representing the SINR $\gamma_{u}$, and the total interference-plus-noise of user $u$, respectively. Constraint (III) is convex, whereas constraint (9a) is nonconvex due to the presence of the norm function on its right-hand side and the product of two variables on the left-hand side. For now, we retain the nonconvex constraint (9a) as it is, since the two-level distributed algorithm is capable of handling it.
@@
$\sum_{\mathcal{U}\setminus\{u\}}\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}}\right|^{2}\leq\sum_{b\in\mathcal{B}}\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|^{2}},~{}\forall u.$ (10)
@@
$\beta_{u}\geq\sum_{b\in\mathcal{B}}\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|^{2}}+\sigma^{2}_{n},\,~{}\forall u.$ (11)
To solve the problem through the two-level distributed algorithm, we introduce a few more new slack variables $\mathbf{s}_{b}=[s_{b,1},\dots,s_{b,U}]^{\top}\in{\mathbb{R}}^{U}~{}\text{and}~{}\mathbf{I}_{b}=[I_{b,1},\dots,I_{b,U}]^{\top}\in{\mathbb{R}^{+}}^{U}$ to (13a) replace constraints (9a) and (11) equivalently as
@@
$\beta_{u}\geq\sum_{b\in\mathcal{B}}{I_{b,u}}+\sigma^{2}_{n},\,~{}\forall u,$
where $s_{b,u}$ and $I_{b,u}$ are defined as follows:
$I_{b,u}\geq\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|^{2}},~{}\forall u,~{}\forall b.$
@@
Constraint (III) is convex; however, to reduce computational complexity, it can be reformulated as a series of second-order cone (SOC) constraints using an exponential cone approximation, as shown in (15a[36]. In this formulation, the parameter $\nu$ controls the accuracy of the approximation, and $k^{u}_{i},\ i=1,\dots,\nu+4,\ \forall u$, are newly introduced slack variables.
@@
After aforementioned transformations, problem (II-B) can be reformulated as
@@
where $\boldsymbol{\beta}$, $\boldsymbol{\alpha}$, $\mathbf{S}$, and $\mathbf{I}$ are vectors and matrices that collect the variables $\beta_{u}$, $\alpha_{u}$, $s_{b,u}$, and $I_{b,u},~{}\forall u,~{}\forall b$, respectively. As can be observed, the set of constraints (II-B) and (13a) are defined per BS, while the set of constraints (12a), (14a), and (15a) are defined per UE. However, these two sets of constraints are coupled through the variables $s_{b,u}$ and $I_{b,u}$. To address this coupling, we introduce new slack variables $\overline{\mathbf{s}}_{b}=[\overline{s}_{b,1},\dots,\overline{s}_{b,U}]^{\top}\in{\mathbb{R}^{+}}^{U}$ and $\overline{\mathbf{I}}_{b}=[\overline{I}_{b,1},\dots,\overline{I}_{b,U}]^{\top}\in{\mathbb{R}^{+}}^{U}$, which are associated with $\mathbf{s}_{b}$ and $\mathbf{I}_{b}$, respectively, such that they satisfy the following equality constraints:
$\mathbf{s}_{b}=\overline{\mathbf{s}}_{b},~{}\mathbf{I}_{b}=\overline{\mathbf{I}}_{b},~{}\forall b.$ (17)
@@
$\overline{I}_{b,u}\geq\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|}^{2},~{}\forall u,~{}\forall b.$
@@
where $\mathcal{C}$ represents a nonconvex constraint set, and $\overline{\mathcal{C}}$ is a compact convex set. These sets are coupled through the equality constraints defined in (17). The problems in the form of (III) can be expressed in the general distributed form as
@@ IV-A Reformulation of Problem (III)
For reformulation, we introduce auxiliary variables $\mathbf{z}_{l,b}\in\mathbb{R}^{U},~{}l\in\mathcal{L}\triangleq\{s,I\},~{}\forall b,$ and replace the equality constraints in (17) with the following constraints as
@@
$\mathbf{I}_{b}-\overline{\mathbf{I}}_{b}+\mathbf{z}_{I,b}=0,\quad\forall b,$
@@
IV-B Inner-Level Three-Block ADMM Algorithm
Given $\boldsymbol{\lambda}^{(k)}_{l,b}$ and $\rho^{(k)}_{o}$ as the Lagrange multipliers and penalty parameter for the $k$th outer-level iteration, the augmented Lagrangian function corresponding to the $k$th inner-level problem is defined in (26).
@@ IV-B Inner-Level Three-Block ADMM Algorithm
In (26) and (27), matrices $\mathbf{Z}_{b},~{}\forall b,$ include the vectors $\mathbf{z}_{l,b},~{}\forall l$, and $\boldsymbol{\psi}_{l,b}$ denotes the Lagrange multipliers associated with the equality constraints in (22a), which are reflected by the defined residuals $\mathbf{r}_{l,b}$, as

### v2

@@ II-B HAPS-UE Channel
$\mathbf{h}_{u}^{B+1}=\frac{1}{\sqrt{{PL}_{(B+1),u}}}\Big(\sqrt{\frac{1}{1+K_{u}}}\overline{\mathbf{h}}_{u}+\sqrt{\frac{K_{u}}{1+K_{u}}}{\color[rgb]{0,0,0}\hat{\mathbf{h}}_{u}}\Big),~\forall u,\qquad\qquad\qquad$ (3)
where ${PL}_{(B+1),u}$ represents the FSPL between HAPS and UE $u$, calculated according to (2). $K_{u}$ is the Rician factor for UE $u$, and $\overline{\mathbf{h}}_{u}\in\mathbb{C}^{N_{B+1}}$ represents the NLoS component of the channel vector with its elements from a normal random distribution with zero mean and unit variance, $\mathcal{NC}(0,1)$. Accordingly, $\color[rgb]{0,0,0}\hat{\mathbf{h}}_{u}$} is the LoS component of the channel vector, given by [35] as
${\color[rgb]{0,0,0}\hat{\mathbf{h}}_{u}}=\mathbf{a}(\theta_{u},\phi_{u})\otimes\mathbf{b}(\theta_{u},\phi_{u}),~\forall u,$ (4)
@@
As mentioned earlier, we adopt the cell-free as the underlying network architecture where all UEs are served by all BSs. We assume that each BS has perfect information regarding the channel coefficients between antenna elements and UEs. Before defining the SINR expression, we first describe the received signal model under the considered cell-free transmission architecture. In this setup, each UE is jointly served by multiple BSs. Following the standard coherent joint-transmission model [36], all serving BSs transmit the same data symbol $s_{u}$ intended for UE $u$, each precoded by a BS-specific beamforming vector $\mathbf{w}^{b}_{u}$. As a result, the desired signal components transmitted from different BSs add coherently at the receiver, while interference arises only from symbols $s_{k}$ intended for other UEs $k\in\mathcal{U}\setminus\{u\}$. The received signal at UE $u$ is therefore given by
$y_{u}=\underbrace{\sum_{b\in\mathcal{B}}(\mathbf{h}^{b}_{u})^{H}\mathbf{w}^{b}_{u}s_{u}}_{\text{combined desired signal}}+\underbrace{\sum_{k\in\mathcal{U}\setminus\{u\}}\sum_{b\in\mathcal{B}}(\mathbf{h}^{b}_{u})^{H}\mathbf{w}^{b}_{k}s_{k}}_{\text{multi-user interference}}+n_{u},~\forall u,\qquad$ (6)
where $(\mathbf{h}^{b}_{u})^{H}$ denotes the channel from BS $b$ to UE $u$, $\mathbf{w}^{b}_{k}\in\mathbb{C}^{N_{b}}$ refers to the beamforming vector for UE $k$ at BS $b$, and $n_{u}$ is the additive white Gaussian noise (AWGN) with variance $\sigma^{2}_{n}$. Since all serving BSs transmit the same symbol $s_{u}$, the UE does not need to decompose per-BS contributions; it simply decodes a single effective combined stream. Based on this received signal model, the SINR of UE $u$ in the cell-free vHetNet can be expressed as [37]
$\gamma_{u}=\cfrac{\left|\sum_{b\in\mathcal{B}}\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{u}\right|^{2}}{\sum_{k\in\mathcal{U}\setminus\{u\}}\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}}\right|^{2}+\sigma^{2}_{n}},~\forall u.$ (7)
Since the PF criterion effectively improves both the minimum SE of the UE and the sum SE [17], we adopt it as the objective function in our study. Accordingly, the PF objective is expressed as the sum of logarithms of SEs over all UEs, where the SE of UE $u$ is defined as $\log_{2}(1+\gamma_{u})$. This objective function follows the classical PF formulation introduced in [38], in which a logarithmic utility function is employed due to its increasing yet concave nature. In this utility, the logarithm naturally promotes fairness by giving more weight to UEs with low SINR and reducing the influence of UEs with already high SINR. Consequently, the performance of weaker UEs improves, while still maintaining strong overall system performance.
Accordingly, the preliminary PFBWD optimization problem can be expressed as
@@
where $P_{b}^{\text{max}}$ is the maximum allowable transmit power of BS $b$, and $\gamma_{\text{min}}$ is the minimum required SINR for each UE. In problem (II-B), constraint (II-B) ensures that each UE $u$ meets its minimum data rate requirement, while constraint (II-B) restricts the total transmit power of each BS $b$ to not exceed $P_{b}^{\text{max}}$. Each BS is subject to its own maximum transmit power constraint, i.e., the HAPS and MBSs are assigned distinct power budgets that reflect their hardware capabilities and link budgets. Problem (II-B) is nonconvex due to the objective function (II-B) and the nonconvex nature of constraint (II-B), rendering it intractable for conventional solvers. Typically, problems of this nature can be addressed by developing an iterative algorithm using the SCA framework. In each iteration of the SCA, a convex approximation of the original problem is solved. The solution obtained at convergence is considered suboptimal relative to the original nonconvex problem. Additionally, the approximated problems can be solved using either a centralized or a distributed algorithm. However, for large-scale problems, it is advisable to avoid a centralized algorithm, as it requires collecting channel coefficients from all BSs to solve the optimization problem, and thereafter distributing the decisions back to the BSs. This process imposes significant signaling overhead on the network. Additionally, the centralized approach must solve a large-scale optimization problem to determine the decision variables for the entire network, resulting in high computational complexity.
The problem at hand is large in scale and complexity, thereby necessitating the development of a distributed algorithm to solve it efficiently. To enable distributed computation, the original centralized formulation presented in (II-B) must first be transformed into an equivalent distributed formulation to which a distributed optimization method can be applied. The key objective of this reformulation is to enable parallel decomposition, whereby the problem is structured in a way that enforces per-BS constraints. This structure allows the overall problem to be divided and solved across multiple BSs simultaneously. In the subsequent section, we focus on this transformation process, aiming to recast the original problem (II-B) into a distributed format suitable for distributed implementation. Although the proposed distributed algorithm is general and applicable to any multi-BS deployment, its advantages become particularly pronounced in cell-free architectures, where each BS serves all UEs simultaneously, resulting in a larger optimization dimension and requiring tight coordination across BSs.
@@ III Distributed Reformulation of the Original Problem (II-B)
In this section, we detail the series of reformulations required to equivalently transform the problem (II-B) to a distributed format that can be solved by the two-level distributed algorithm. To this end, first, we deal with the objective function (II-B). We introduce $U$ slack variables $t_{u},~\forall u,$ as the lower bound for $\log_{2}(1+\gamma_{u}),~\forall u$. Accordingly, the original objective function (II-B) can be equivalently replaced by $f(\mathbf{t})=\sum_{u=1}^{U}\log_{2}(t_{u})$, where $\mathbf{t}=[1,\dots,t_{U}]^{\top}\in{\mathbb{R}^{+}}^{U}$. This reformulation introduces the following $U$ additional constraints:
@@
Next, to decouple the interference and desired signal terms in $\gamma_{u}$, we define slack variables $\alpha_{u},~\forall u,$ and $\beta_{u},~\forall u,$ representing the SINR $\gamma_{u}$, and the total interference-plus-noise of UE $u$, respectively. Accordingly, the two constraints below will be defined as
@@
$\beta_{u}\geq\sum_{k\in\mathcal{U}\setminus\{u\}}\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}}\right|^{2}+\sigma^{2}_{n},\,~\forall u.$
Constraint (III) is convex, whereas constraint (10a) is nonconvex due to the presence of the norm function on its right-hand side and the product of two variables on the left-hand side. For now, we retain the nonconvex constraint (10a) as it is, since the two-level distributed algorithm is capable of handling it.
@@
$\sum_{k\in\mathcal{U}\setminus\{u\}}\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}}\right|^{2}\leq\sum_{b\in\mathcal{B}}\sum_{k\in\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|^{2}},~\forall u.$ (11)
@@
$\beta_{u}\geq\sum_{b\in\mathcal{B}}\sum_{k\in\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|^{2}}+\sigma^{2}_{n},\,~\forall u.$ (12)
To solve the problem through the two-level distributed algorithm, we introduce a few more new slack variables $\mathbf{s}_{b}=[s_{b,1},\dots,s_{b,U}]^{\top}\in{\mathbb{R}}^{U}~\text{and}~\mathbf{I}_{b}=[I_{b,1},\dots,I_{b,U}]^{\top}\in{\mathbb{R}^{+}}^{U}$, where $s_{b,u}$ and $I_{b,u}$ are defined as follows:
$I_{b,u}\geq\sum_{k\in\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|^{2}},~\forall u,~\forall b.$
Consequently, constraints (10a) and (12) will equivalently be replaced as
@@
$\beta_{u}\geq\sum_{b\in\mathcal{B}}{I_{b,u}}+\sigma^{2}_{n},\,~\forall u.$
@@
As demonstrated in Appendix A, all inequality constraints introduced in the reformulation are active at optimality, ensuring that the reformulated problem is tight and equivalent to the original formulation. Constraint (III) is convex; however, to reduce computational complexity, it can be reformulated as a series of second-order cone (SOC) constraints using an exponential cone approximation, as shown in (16a[39]. In this formulation, the parameter $\nu$ controls the accuracy of the approximation, and $k^{u}_{i},\ i=1,\dots,\nu+4,\ \forall u$, are newly introduced slack variables.
@@
After aforementioned transformations, problem (II-B) can be equivalently reformulated as
@@
where $\boldsymbol{\beta}$, $\boldsymbol{\alpha}$, $\mathbf{S}$, and $\mathbf{I}$ are vectors and matrices that collect the variables $\beta_{u}$, $\alpha_{u}$, $s_{b,u}$, and $I_{b,u},~\forall u,~\forall b$, respectively. As can be observed, the set of constraints (II-B) and (13a) are defined per BS, while the set of constraints (14a) and (15a) are defined per UE. However, these two sets of constraints are coupled through the variables $s_{b,u}$ and $I_{b,u}$. To address this coupling and facilitate the two-level distributed algorithm, we introduce new slack variables $\overline{\mathbf{s}}_{b}=[\overline{s}_{b,1},\dots,\overline{s}_{b,U}]^{\top}\in{\mathbb{R}^{+}}^{U}$ and $\overline{\mathbf{I}}_{b}=[\overline{I}_{b,1},\dots,\overline{I}_{b,U}]^{\top}\in{\mathbb{R}^{+}}^{U}$, which are associated with $\mathbf{s}_{b}$ and $\mathbf{I}_{b}$, respectively, such that they satisfy the following equality constraints:
$\mathbf{s}_{b}=\overline{\mathbf{s}}_{b},~\forall b,$ (18a)
$\mathbf{I}_{b}=\overline{\mathbf{I}}_{b},~\forall b.$
@@
$\overline{I}_{b,u}\geq\sum_{k\in\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|}^{2},~\forall u,~\forall b.$
@@
where $\mathcal{C}$ represents a nonconvex constraint set, and $\overline{\mathcal{C}}$ is a compact convex set. These sets are coupled through the equality constraints defined in (18a). The problems in the form of (III) can be expressed in the general distributed form as
@@ IV-A Reformulation of Problem (III)
For reformulation, we introduce auxiliary variables $\mathbf{z}_{l,b}\in\mathbb{R}^{U},~l\in\mathcal{L}\triangleq\{s,I\},~\forall b,$ and replace the equality constraints in (18a) with the following constraints as
@@
$\mathbf{I}_{b}-\overline{\mathbf{I}}_{b}+\mathbf{z}_{I,b}=0,\quad\forall b.$
@@
IV-B Inner-Level Three-Block ADMM Algorithm
Given $\boldsymbol{\lambda}^{(k)}_{l,b}$ and $\rho^{(k)}_{o}$ as the Lagrange multipliers and penalty parameter for the $k$th outer-level iteration, the augmented Lagrangian function corresponding to the $k$th inner-level problem is defined in (27). In (27) and (28), matrices $\mathbf{Z}_{b},~\forall b,$ include the vectors $\mathbf{z}_{l,b},~\forall l$, and $\boldsymbol{\psi}_{l,b}$ denotes the Lagrange multipliers associated with the equality constraints in (23a), which are reflected by the defined residuals $\mathbf{r}_{l,b}$, as